OpenStudy (lyssakat):

Will fan and medal! A circle is described by the equation (x−1/2)^2+(y−3/2)^2=4/9. What are the cordinates for the center of the circle and the length of the radius?

OpenStudy (faiqraees):

\[\large\rm (x- \color{#0000ff}{ \displaystyle a})^2+(y-\color{#0000ff}{ \displaystyle b})^2=\color{#0000ff}{ \displaystyle r}^2\] where (a,b) are the x and y coordinates of the center of the cirle and r is the radius of the circle

OpenStudy (lyssakat):

ok... so i just plug in the numbers with the equation??

OpenStudy (faiqraees):

Just find out the value of a and b

OpenStudy (lyssakat):

its 1/2 and 3/2 ?

OpenStudy (phi):

yes

OpenStudy (faiqraees):

yes

OpenStudy (lyssakat):

oh okay what about the units?

OpenStudy (faiqraees):

no units

OpenStudy (lyssakat):

Well it has answer choices with different units. (-1/2,-3/2), 23units (-1/2,-3/2), 49units (1/2,3/2), 49units (1/2,3/2), 23units it asks for units

OpenStudy (faiqraees):

oh units is for radius

OpenStudy (lyssakat):

Oh okay!!! thank you!!

OpenStudy (anonymous):

(a, b)=(12 , 32 ), r=​23

OpenStudy (lyssakat):

Oh dang it was wrong :(

OpenStudy (lyssakat):

it was 23

OpenStudy (anonymous):

(x−12 )2+(y−32 )2=49 (x−a)2+(y−b)2=r2 is the circle equation with a radius r, centered at (a, b) $\left(x-\frac{1}{2}\right)^2+\left(y-\frac{3}{2}\right)^2=\left(\frac{2}{3}\right)^2$(x−12 )2+(y−32 )2=(23 )2 (a, b)=(12 , 32 ), r=​23

OpenStudy (anonymous):

(a, b)=(1/2 , 3/2 )​, r=23

OpenStudy (anonymous):

no its not 23 dear is 2/3